OpenStudy (amilapsn):

For this christmas i m gonna make a christmas tree. I want to know the length of the wire needed to wrap around it. Christmas tree can be considered as a cone of heigt H and radius R. The gap between wire should be h in vertical direction.

OpenStudy (amilapsn):

⬆⬆sketch🌲🌲

OpenStudy (mhchen):

I think this has to do with transformations of a sine/cosine function. Unfortunately, I'm supposed to learn that next semester .-.

OpenStudy (amilapsn):

My brother says that we can get a approximation like \[\frac {H}{h}\times 2 \pi R\]

OpenStudy (mhchen):

mmm,, that would be a good approximation if the tree was a cylinder. I'm gonna keep looking into polar equations.

OpenStudy (amilapsn):

But I dont know how it works

OpenStudy (mhchen):

Yeah, me neither xD, but I think I got somewhere. Check this out: https://www.desmos.com/calculator I'm sure you're familiar with this graphing site. In the equation box on the left, type in: r = -1.1^theta

OpenStudy (amilapsn):

Would it correct if we get the half of my brother's approximation?

OpenStudy (mhchen):

It would be okay, but not necessarily exact. We have to use integrals for this. Tag a calculus person, they should know this.

OpenStudy (amilapsn):

I did this for H=12 R=3 and h=.5

OpenStudy (amilapsn):

Approximation givea almost the same answer

OpenStudy (mhchen):

WOah....okay that's beyond the scope of what I know. This question actually makes me wonder really. Like if i were to think about it logically:  I'm trying to think of what a third rotation wuold be like but aaah, my brain.